Write an equation in slop-intercept form for the line with the given condition that contains the point (6,-2): Parallel to the line 2x+y=5
Answer by venugopalramana(3286)

Write an equation in slop-intercept form for the line with the given condition that contains the point (6,-2):
Parallel to the line 2x+y=5
EQN. OF PARALLEL LINE WILL BE
2X+Y =K
IT PASSES THROUGH (6,-2).HENCE
2*6-2=K
K=10
EQN. IS
2X+Y=10
Y=-2X+10
SLOPE = M = -2
INTERCEPT = 10